# Hugo Malard1  Michel Olvera1  Stéphane Lathuiliere1  Slim Essid1 1LTCI, Télécom Paris, Institut Polytechnique de Paris

### v1

@@ Abstract
Large-scale pre-trained audio and image models demonstrate an unprecedented degree of generalization, making them suitable for a wide range of applications. Here, we tackle the specific task of sound-prompted segmentation, aiming to segment image regions corresponding to objects heard in an audio signal. Most existing approaches tackle this problem by fine-tuning pre-trained models or by training additional modules specifically for the task. We adopt a different strategy: we introduce a training-free approach that leverages Non-negative Matrix Factorization (NMF) to co-factorize audio and visual features from pre-trained models to reveal shared interpretable concepts. These concepts are passed to an open-vocabulary segmentation model for precise segmentation maps. By using frozen pre-trained models, our method achieves high generalization and establishes state-of-the-art performance in unsupervised sound-prompted segmentation, significantly surpassing previous unsupervised methods.
@@ 1 Introduction
Audio-visual perception has attracted considerable interest due to its practical applications in various fields, especially robotics, video understanding [4, 39], and recently audio-visual language modeling [25, 6]. A particularly compelling task in this area is sound-prompted segmentation, which consists in pinpointing the image regions associated with sounds in a corresponding audio signal. Most approaches proceed to align audio and visual representations through contrastive learning techniques [41, 38, 30]. These techniques leverage paired audio-visual data to learn a shared embedding space, where corresponding audio and visual features are brought closer together, while non-corresponding pairs are pushed apart. The main challenge lies in learning local features that can be used for segmenting audio-related concepts, as videos typically provide global-level alignment (as opposed to local alignment where each part of the image is matched with a part of the audio). More recently, a few works have made significant progress in this direction through the use of pre-trained models. For example, Park et al. [31] use a pre-trained CLIP encoder [32] and audio tokenizer within a relatively complex framework that translates audio signals into tokens compatible with CLIP’s text encoder. This is combined with a grounding mechanism, enabling audio-driven segmentation of the image. Similarly, Hamilton et al.’s DenseAV  [17] identifies image regions linked to sounds without supervision. Using a DINO image backbone [1] and the HUBERT audio transformer [19], it employs multi-head feature aggregation for contrastive learning on video-audio pairs.
@@
While these approaches tackle sound-prompted segmentation through carefully designed architectures and dedicated training procedures, we propose a radically different strategy. Specifically, we ask: can this task be addressed using only pre-trained audio and image models, without any task-specific training, by directly analyzing their feature representations?
To answer this question, we introduce TACO: a Training-free Audio-visual CO-factorization approach, based on soft Non-negative Matrix co-Factorization (soft co-NMF) which identifies audio and visual tokens that are co-activated in the input signal. By leveraging pre-trained CLIP and CLAP [10] backbones for image and audio, our co-NMF formulation enables audio-visual correspondence analysis without the need for training. Specifically, using a set of semantic anchor words projected into the CLIP and CLAP spaces, we are able to match audio features with visual ones in an interpretable fashion. Through simple inference-time optimization, our framework segments both visual and audio signals, performing object segmentation and sound localization in a training-free manner (illustrated in Figure 1). Furthermore, the interpretable factors identified by our decomposition framework allow us to directly prompt a pre-trained open-vocabulary segmenter, such as FC-CLIP [47], enhancing segmentation quality while remaining zero-shot.
Our approach stands out for its training-free paradigm and the interpretability offered by the NMF framework. Additionally, leveraging the robustness of the frozen CLIP and CLAP models enables us to achieve significantly superior performance compared to existing unsupervised training-based methods on established benchmarks. To summarize, our contributions are:
  • We introduce TACO, the first training-free approach for sound-prompted segmentation based on pre-trained deep representations. Our framework leverages the inherent interpretability of NMF, enabling clear, semantically grounded visualization of interactions between segmentation outputs and the concepts identified within the signal.
  • We enhance the quality of the segmentation provided by our co-NMF framework by incorporating a pre-trained open-vocabulary model (namely, FC-CLIP [47]), prompting its decoder with the most activated factors identified by our NMF system, for improved accuracy.
  • We rigorously validate our approach through extensive experiments, providing both quantitative and qualitative analyses. We demonstrate the superior performance of our training-free approach across four datasets and multiple variants of the sound-prompted segmentation task. Additionally, the code used in our experiments will be released.
@@ 2 Related Work
2.1 Sound-Prompted Segmentation
@@
Sound-prompted segmentation is the task of localizing in an image the object emitting the sound that is heard on an associated audio. Recent methods primarily utilize cross-modal attention for sound-prompted segmentation [42, 36, 35], often combined with a contrastive loss that learns to align audio and image global representations, hypothesizing that this process would implicitly foster the emergence of local alignment. Extending this contrastive learning framework, numerous enhancements have been introduced. These include integrating challenging negatives from background areas [3], adopting iterative contrastive learning with pseudo-labels from earlier model epochs [22], maintaining transformation invariance and equivariance through geometric consistency [23], leveraging semantically similar hard positives [37], and implementing false negative-aware contrastive learning through intra-modal similarities [40].
@@
Very recently, large pre-trained models have started to be used in order to benefit from their rich data representation. ACL-SSL [31] relies on the CLIP image encoder as well as a segmentation model CLIP-Seg[24], and learns an additional module that projects the audio in the CLIP space. Relying on DINO and HUBERT (further fine-tuned), Hamilton et al.[17] trained a multi-head module that identifies the localization of the sound in the image features.
Unlike these methods, our framework relies on a training-free setup leveraging pre-trained models and matrix decomposition. By avoiding training, our method not only avoids the need for a training dataset but also preserves the generalization abilities of the pre-trained model.
2.2 NMF and Deep Feature Factorization
@@
Non-negative matrix factorization (NMF) has found applications in diverse fields, including audio source separation [15], document clustering [44], or face recognition [16], to mention a few. Previous research has expanded NMF to multiple layers [7], implemented NMF using neural networks [9], and utilized NMF approximations as inputs to neural networks[43]. More recently it has been applied to decompose deep neural features, for concept discovery [8], interpretability [27, 28] or audio-visual source separation [29]. While it has already been applied to co-factorize audio and visual (handcrafted) features [34], to the best of our knowledge this work is the first one to apply it to deep features for sound-prompted segmentation, as well as the first attempt to co-factorize representations from different spaces while enforcing semantic matching.
@@
3 [Uncaptioned image] TACO: Training-free Audio-visual CO-factorization
This section introduces our method, TACO, which allows training-free sound-prompted segmentation. Section 3.1 starts by presenting the main motivations. Section 3.2 provides the necessary background on NMF, and section 3.3 introduces our framework for extracting audio-visual concepts using NMF, while section 3.4 describes our complete solution for audio-source segmentation, which integrates our factorization approach into a pre-trained open-vocabulary image segmenter, enabling fine segmentation in a training-free manner.
3.1 Motivation
@@
Our methodology leverages the capacity of recent large pre-trained models to effectively localize patterns within their input signals [47]. Localization here refers to identifying spatial regions in images and temporal segments in audio. From this observation, we hypothesize that separately trained audio and visual encoders contain all the necessary information for sound-prompted segmentation without requiring additional training. To achieve this, we propose a method, TACO, which identifies concepts that co-activate across audio and image representations exploiting the Non-negative Matrix Factorization (NMF) framework [21]. NMF provides a training-free approach that jointly decomposes each signal into an activation matrix and a concept matrix. The activation matrix, in particular, can be interpreted as a segmentation matrix, highlighting regions of concept alignment.
@@
Before detailing our approach, it is important to note that NMF is a tool that is meant to interpret input matrices composed of only non-negative values. The non-negativity constraints are key to the interpretability offered by the NMF framework. However, deep representations often include negative coefficients. To address this, a straightforward solution is to clip all negative values to zero, thus ensuring matrix non-negativity. Although this adjustment may result in some information loss, our preliminary experiments indicate that with backbones like CLIP [32] and CLAP [10], the representation power remains unaffected for our specific tasks and settings. Further experimental details on this are provided in Sec. 4.
@@ 3.2 Background
Non-negative Matrix Factorization (NMF) [21] involves decomposing a non-negative matrix $X\in\mathbb{R}^{N\times C}_{+}$ into two non-negative matrices $U\in\mathbb{R}^{N\times K}_{+}$ and $V\in\mathbb{R}^{K\times C}_{+}$ such that $X\approx UV$. In typical machine learning applications of NMF [12], $N$ represents the number of observations and $C$ denotes the number of (non-negative) features per observation, while $K$ is the rank of the factorization, controlling the number of factors or components into which the original matrix is decomposed. Assuming a measure of fit $D$ is considered, NMF can be formulated as the following constrained optimization problem:
@@
As the $V$ matrix is interpreted as the factor matrix, $U$ corresponds to the activations of those factors.
@@
extends the NMF framework to multiple data views analyzed in parallel. Assuming two data matrices $X_{1}\in\mathbb{R}^{N\times C}_{+}$ and $X_{2}\in\mathbb{R}^{N\times C}_{+}$ containing different modalities (such as audio and visual streams of a video), co-NMF exploits the mutual information shared between these modalities. It assumes that the different modalities share a common activation matrix $U$. In particular, [45, 46] decomposes these matrices into a shared activation matrix $U\in\mathbb{R}^{K\times C}_{+}$ with view-specific factors $V_{1}\in\mathbb{R}^{N\times K}_{+}$ and $V_{2}\in\mathbb{R}^{N\times K}_{+}$. This formulation forces the activation matrix $U$ to be consistent across both modalities, reflecting the intuition that events occurring in the audio also happen in the same time in the image stream. Although the factors are encoded in different spaces, they should remain the same. Formally, assuming measures $D_{1}$ and $D_{2}$, and weighting hyperparameters $\beta_{1}>0$, $\beta_{2}>0$, co-NMF solves the following problem:
$\min_{U,V_{1},V_{2}}\beta_{1}D_{1}(X_{1}|UV_{1})+\beta_{2}D_{2}(X_{2}|UV_{2}),$ (2)
@@
Note that, as there might be some temporal shift between the audio and image events, enforcing the same $U$ matrix for both decompositions can be overly restrictive. To mitigate this issue, soft non-negative Matrix Co-Factorization (soft co-NMF) [34] introduces two separate activation matrices and includes a penalty term to minimize their dissimilarity. It is obtained by solving the following optimization problem, with a penalty function $P$, and hyperparameter $\beta_{p}>0$
$\min_{U_{1},U_{1},V_{2},V_{2}}\beta_{1}D_{1}(X_{1}|U_{1}V_{1})+\beta_{2}D_{2}(X_{2}|U_{2}V_{2})$ (3)
$+\beta_{p}P(U_{1},V_{1},U_{2},V_{2}),$
@@
3.3 Deep Non-negative Matrix co-Factorization
This section details how TACO leverages soft co-NMF to perform a zero-shot sound-prompted segmentation.

### v3

@@ Abstract
Large-scale pre-trained audio and image models demonstrate an unprecedented degree of generalization, making them suitable for a wide range of applications. Here, we tackle the specific task of sound-prompted segmentation, aiming to segment image regions corresponding to objects heard in an audio signal. Most existing approaches tackle this problem by fine-tuning pre-trained models or by training additional modules specifically for the task. We adopt a different strategy: we introduce a training-free approach that leverages Non-negative Matrix Factorization (NMF) to co-factorize audio and visual features from pre-trained models so as to reveal shared interpretable concepts. These concepts are passed on to an open-vocabulary segmentation model for precise segmentation maps. By using frozen pre-trained models, our method achieves high generalization and establishes state-of-the-art performance in unsupervised sound-prompted segmentation, significantly surpassing previous unsupervised methods.
@@ 1 Introduction
Audio-visual perception has attracted considerable interest due to its practical applications in various fields, especially robotics, video understanding [4, 42], and recently, audio-visual language modeling [28, 6]. A particularly compelling task in this area is sound-prompted segmentation [18]—also known as audio-visual segmentation [53] —which involves identifying image regions associated with sounds in the accompanying audio. Whereas several prior approaches frame this task as a supervised segmentation problem [25, 11, 24], some explore unsupervised learning, relying solely on inherent cross-modal features rather than ground-truth audiovisual masks. Most of them proceed to aligning audio and visual representations through contrastive learning techniques [43, 41, 33]. Such techniques exploit paired audio-visual data to learn a shared embedding space, where corresponding audio and visual features are brought closer together, while non-corresponding pairs are pushed apart. The main challenge lies in learning local features that can be used for segmenting audio-related concepts, as videos typically provide global-level alignment (as opposed to local alignment, where each part of the image has a match with a defined part of the audio). Recently, a few works have made significant progress in this direction through the use of pre-trained models. Park et al. [34] use a pre-trained CLIP encoder [35] and audio tokenizer within a relatively complex framework that translates audio signals into tokens compatible with CLIP’s text encoder. This is combined with a grounding mechanism, enabling audio-driven segmentation of the image. Similarly, Hamilton et al.’s DenseAV  [18] identifies image regions linked to sounds without supervision. Using a DINO image backbone [1] and the HUBERT audio transformer [19], it employs multi-head feature aggregation for contrastive learning on video-audio pairs.
These approaches tackle sound-prompted segmentation through carefully designed architectures and dedicated training procedures, which often compromise the model’s generalization capabilities and abilities in other tasks. Here, we propose a radically different strategy.
@@
Specifically, we ask: Can sound-prompted visual segmentation be achieved using only pre-trained audio and image models, in an unsupervised manner, by analyzing their frozen feature representations?
To answer this question, we introduce TACO: a Training-free Audio-visual CO-factorization approach, which identifies audio and visual tokens that are co-activated in the input signal. Leveraging pre-trained CLIP and CLAP [10] backbones for image and audio, respectively, we introduce a semantically constrained Non-negative Matrix co-Factorization (Sem co-NMF) framework. This co-NMF formulation enables audio-visual correspondence analysis without the need for training. Specifically, using a set of semantic anchor words projected into the CLIP and CLAP spaces, we are able to match audio features with visual ones in an interpretable fashion. Through simple inference-time optimization, our framework establishes local correspondence between visual and audio signals, and further performs sound-prompted segmentation in a training-free manner, as illustrated in Figure 1. Furthermore, the interpretable factors identified by our decomposition framework allow us to directly prompt a pre-trained open-vocabulary segmenter, such as FC-CLIP [50], enhancing segmentation quality while remaining zero-shot.
Our unsupervised approach stands out for its training-free paradigm which preserves the generalization capabilities of the pretrained models, as well as the interpretability offered by the NMF framework. This training-free setup holds great potential for multi-task scenarios, enabling a single model to support multiple downstream applications without additional fine-tuning. Additionally, leveraging the robustness of the frozen CLIP and CLAP models enables us to achieve significantly superior performance compared to existing unsupervised training-based methods on established benchmarks. To summarize, our contributions are:
  • We introduce TACO, the first training-free approach for sound-prompted segmentation based on pre-trained deep representations. Our framework leverages the inherent interpretability of matrix factorization, enabling clear, semantically grounded visualization of interactions between segmentation outputs and the concepts identified within the signal.
  • We revisit soft co-NMF for audiovisual sound source localization by constraining the semantics of the decomposition and incorporating a novel penalty function. This new decomposition (called Sem co-NMF) enables finding audio-visual correspondences across unaligned sound and image representation backbones.
  • Our Sem co-NMF framework is meant to extract both an audiovisual segmentation and encode audiovisual concepts that can be incorporated into a pre-trained open-vocabulary model (namely, FC-CLIP [50]), prompting its decoder with the most activated factors identified by our decomposition system, allowing significant leap in accuracy.
  • We rigorously validate our approach through extensive experiments, providing both quantitative and qualitative analyses. We demonstrate the superior performance of our unsupervised (training-free) approach across four datasets and multiple variants of the sound-prompted segmentation task. The code will be released as open-source upon acceptance.
@@ 2 Related Work
Unsupervised Sound-Prompted Segmentation
@@
The sound-prompted segmentation task aims to localize the object in an image that emits the sound present in the associated audio. While often referred to as audio-visual segmentation [53] , we adopt the term used in [18] throughout this paper, as it more clearly indicates that only the image is segmented, not the audio. Recent methods primarily utilize cross-modal attention for sound-prompted segmentation [45, 39, 38], often combined with a contrastive loss aligns audio and image global representations, hypothesizing that this process would implicitly foster the emergence of local alignment. Extending this contrastive learning framework, numerous enhancements have been introduced. These include integrating challenging negatives from background areas [2], adopting iterative contrastive learning with pseudo-labels from earlier model epochs [23], maintaining transformation invariance and equivariance through geometric consistency [26], leveraging semantically similar hard positives [40], and implementing false negative-aware contrastive learning via intra-modal similarities [44].
@@
More recently, large pre-trained models have started to be used in order to benefit from their rich data representation. ACL-SSL [34] relies on the CLIP image encoder as well as a segmentation model CLIP-Seg [27], and learns an additional module that projects the audio in the CLIP space. Relying on DINO and HUBERT (further fine-tuned), Hamilton et al. [18] trained a multi-head module that identifies the localization of the sound in the image features. Unlike these methods, our framework uses a training-free setup with pre-trained models and matrix decomposition, requiring no training data and preserving the model’s generalization abilities.
NMF and Deep Feature Factorization
@@
Non-negative matrix factorization (NMF) has been applied in diverse fields, including audio source separation [16], document clustering [47], and face recognition [17]. Previous research has expanded NMF to multiple layers [7], implemented NMF using neural networks [9], and utilized NMF approximations as inputs to neural networks[46]. More recently it has been applied to decompose deep neural features, for concept discovery [8], interpretability [31, 30] or audio-visual source separation [32]. While it has already been applied to co-factorize audio and visual (handcrafted) features [37], to the best of our knowledge, this is the first work applying it to deep features for sound-prompted segmentation, and the first to co-factorize representations from different spaces while enforcing semantic matching.
@@
3 [Uncaptioned image] TACO: Training-free semantically constrained Audio-visual CO-factorization
@@
Our methodology leverages the capacity of recent large pre-trained models to effectively localize patterns within their input signals [50]. Localization here refers to identifying spatial regions in images and spectro-temporal segments in audio. From this observation, we hypothesize that separately trained audio and visual encoders contain all the necessary information for sound-prompted segmentation without requiring additional training. To achieve this, we propose a method, TACO, which identifies concepts that co-activate across audio and image representations exploiting the Non-negative Matrix Factorization (NMF) framework [22]. NMF provides a training-free approach that jointly decomposes each signal into an activation matrix and a concept matrix. The activation matrix, in particular, can be interpreted as a segmentation matrix, highlighting regions of concept alignment.
@@
Before detailing our approach, it is important to note that NMF is a tool that is meant to interpret input matrices composed of only non-negative values. The non-negativity constraints are key to the interpretability offered by the NMF framework. However, deep representations often include negative coefficients. A simple solution is to clip these negative values to zero, ensuring matrix non-negativity. While this may lead to some information loss, our preliminary experiments show that with backbones like CLIP [35] and CLAP [10], representation power remains unaffected for our tasks (further details are in Sec. 4). Section 3.1 provides the necessary background on NMF, and section 3.2 introduces our framework for extracting audio-visual concepts using NMF, while section 3.3 describes our complete solution for audio-source segmentation, which integrates our factorization approach into a pre-trained open-vocabulary image segmenter, enabling fine segmentation in a training-free manner.
@@ 3.1 Background
Non-negative Matrix Factorization (NMF) [22] involves decomposing a non-negative matrix $X\in\mathbb{R}^{N\times C}_{+}$ into two non-negative matrices $U\in\mathbb{R}^{N\times K}_{+}$ and $V\in\mathbb{R}^{K\times C}_{+}$ such that $X\approx UV$. In typical machine learning applications of NMF [13], $N$ represents the number of observations and $C$ denotes the number of (non-negative) features per observation, while $K$ is the rank of the factorization, controlling the number of factors or components into which the original matrix is decomposed. Given a distance $D$ in the matrix space, NMF can be formulated as the following constrained optimization problem:
@@
The matrix $V$ is called factor matrix, and the $U$ matrix corresponds to the activations of those factors.
@@
extends the NMF framework to multiple data views analyzed in parallel. Assuming two data matrices $X_{1}\in\mathbb{R}^{N\times C}_{+}$ and $X_{2}\in\mathbb{R}^{N\times C}_{+}$ containing different modalities (such as audio and visual streams of a video), co-NMF exploits the mutual information shared between these modalities. It assumes that the different modalities share a common activation matrix $U$. In particular, [48, 49] decomposes these matrices into a shared activation matrix $U\in\mathbb{R}^{K\times C}_{+}$ with view-specific factors $V_{1}\in\mathbb{R}^{N\times K}_{+}$ and $V_{2}\in\mathbb{R}^{N\times K}_{+}$. This formulation forces the activation matrix $U$ to be consistent across both modalities, reflecting the intuition that events occurring in the audio also happen at the same time in the image stream. Although the factors are encoded in different spaces, they should remain the same. Formally, given distances $D_{1}$ and $D_{2}$, co-NMF solves the following problem:
$\min_{U,V_{1},V_{2}}D_{1}(X_{1}|UV_{1})+D_{2}(X_{2}|UV_{2}),$ (2)
@@
Note that, as there might be some small temporal shift between the audio and image events, enforcing the same $U$ matrix for both decompositions can be overly restrictive. To mitigate this issue, soft non-negative Matrix Co-Factorization (soft co-NMF) [37] introduces two separate activation matrices and includes a penalty term to minimize their dissimilarity. They are obtained by solving the following optimization problem, with a penalty function $P$, and hyperparameter $\beta_{p}>0$
$\min_{U_{1},U_{1},V_{2},V_{2}}D_{1}(X_{1}|U_{1}V_{1})+D_{2}(X_{2}|U_{2}V_{2})+\beta_{p}P(U_{1},V_{1},U_{2},V_{2}),$ (3)
@@
3.2 Sem Co-NMF: Revisiting soft co-NMF
This section details our redefinition of soft co-NMF to semantically constrained co-NMF (Sem Co-NMF) and its application in TACO.
Soft co-NMF for token decomposition
We assume access to an input video with a corresponding audio signal. For simplicity, let us initially consider the central frame of the video; we will later extend our approach to handle the entire video, incorporating the temporal dimension. Additionally, we assume the availability of trained audio and image backbones. Feeding the audio signal into the audio backbone produces a matrix $X_{A}\in\mathbb{R}^{N_{T}\times C_{A}}_{+}$, where $N_{T}$ is the number of tokens, and $C_{A}$ is the token dimension. Similarly, the image encoder maps an image to a feature representation $X_{I}\in\mathbb{R}^{HW\times C_{I}}_{+}$, where $H$ and $W$ represent the spatial dimensions, and $C_{I}$ is the number of channels. Note that the spatial dimensions are flattened, resulting in a 2D matrix compatible with NMF inputs.
In the original formulation, soft co-NMF factorizes visual and audio matrices containing temporal features (i.e. the first dimension corresponds to time). The penalty term is applied to the activations to ensure that concepts align temporally in both audio and image. However, in our case, the first dimension of the image feature corresponds to space (as we aim for spatial decomposition for segmentation), hence the revision of the factorization model.
In our Sem co-NMF scheme, we aim at factorizing $X_{A}$ into matrices $U_{A}\in\mathbb{R}^{N_{T}\times K}_{+}$ and $V_{A}\in\mathbb{R}^{K\times C_{A}}_{+}$, along with $X_{I}$ into $U_{I}\in\mathbb{R}^{HW\times K}_{+}$ and $V_{I}\in\mathbb{R}^{K\times C_{I}}_{+}$, where $K$ denotes the predefined number of factors. Restricting the decomposition of the $U_{I}$ and $U_{A}$ matrices to the interval $[0,1]$ instead of $\mathbb{R}_{+}$ provides an interpretable, well-calibrated scale, where each concept is either activated (near one) or inactive (near zero) in each token. Thus, we re-parameterize the problem using the sigmoid function $\sigma$ and optimize matrices $\tilde{U}_{A}\in\mathbb{R}^{N_{T}\times K}$ and $\tilde{U}_{I}\in\mathbb{R}^{HW\times K}$, such that $U_{A}=\sigma(\tilde{U}_{A})$ and $U_{I}=\sigma(\tilde{U}_{I})$ with $U_{I}\in[0,1]^{HW\times K}$ and $U_{A}\in[0,1]^{N_{T}\times K}$. Specifically, the $k$-th column of $U_{I}$ and $U_{A}$ can be interpreted as the activation of the $k^{th}$ factor for each token. In this way, the Hadamard product between the original tokens and a column of $U_{I}$ or $U_{A}$ can be viewed as soft masking, highlighting the portion of the original input that activates the $k^{th}$ factor. To jointly decompose the image and audio features, we cannot constrain the activations to be identical, as images encode spatial information and audio captures spectro-temporal patterns. Additionally, constraining the factors $V_{A}$ and $V_{I}$ is challenging, as they reside in different spaces. To address this, we introduce a penalty function that ensures semantic proximity between the decompositions.
Penalty function
Applying soft co-NMF to a specific problem requires careful definition of the penalty term $P$ in Eq. (3), as it impacts both the quality of the factorization and its interpretability. To achieve a semantically consistent factorization of audio-visual observations, $P$ must effectively capture the semantic similarity between the image and audio features. However, since the backbones (e.g., CLIP and CLAP) are not aligned, audio and image factors are encoded in two different subspaces, making $L_{1}$ or $L_{2}$ norms unsuitable as penalty functions. Moreover, audio and image representations are not necessarily commensurable: they often differ in dimensionality, precluding direct use of standard distances. To address this, we propose projecting the audio and visual feature representations into a unified feature space to enable direct comparison. Specifically, we estimate aligned semantic descriptors $\mathcal{D}_{I^{k}}$ and $\mathcal{D}_{A^{k}}$ for the $k^{th}$ column of $U_{I}$ and $U_{A}$, respectively. These descriptors are computed using $J$ semantic anchors, $\{b_{I}^{j}\}_{j=1}^{J}$ and $\{b_{A}^{j}\}_{j=1}^{J}$, which lie in the same spaces as $X_{I}$ and $X_{A}$. Each anchor pair $(b_{I}^{j},b_{A}^{j})$ represents a shared semantic concept, enabling the estimation of aligned semantic descriptors. Specifically, to estimate aligned semantic anchors, we utilize CLIP [35] and CLAP [10] as image and audio backbones. Since these encoders are aligned with text encoders, we can exploit their respective text embeddings to align identical words across audio and image spaces. Thus, we define $b_{I}^{j}=\mathcal{E}_{I}(t_{j})$ and $b_{A}^{j}=\mathcal{E}_{A}(t_{j})$, where $\mathcal{E}_{I}$ and $\mathcal{E}_{A}$ are the text encoders of CLIP and CLAP, respectively, and $\{t_{j}\}_{j=1}^{J}$ represents a predefined word bank.